# MotionCtrl: A Unified and Flexible Motion Controller for Video Generation https://wzhouxiff.github.io/projects/MotionCtrl/

### v1

@@ 4.3 Ablation Studies
Table 3: Ablation of Object Motion Control. The Object Motion Control Module (OMCM), when initially trained on dense object movement trajectories and subsequently fine-tuned with sparse trajectories, outperforms versions trained exclusively on either dense or sparse trajectories.
Training Strategy. Given the limitations of the available training dataset, we propose a multi-step training strategy for MotionCtrl, starting with the Camera Motion Control Module (CMCM) using Realestate10K [33], followed by the Object Motion Control Module (OMCM) with synthesized object movement trajectories. To thoroughly assess our approach, we experiment with reversing the order, training OMCM before CMCM. This sequence does not impact camera motion control, as OMCM components do not participate in CMCM training. However, it does lead to a decrease in object motion control performance. The subsequent training of CMCM adjusts parts of LVDM’s temporal transformers, disrupting the object motion control adaptation achieved during OMCM’s initial training (evidenced by a higher ObjCM score of $25.2712$ compared to $25.1198$ attained with our proposed training sequence). Thus, our multi-step strategy, though a compromise due to dataset constraints, is deliberately structured to train CMCM before OMCM, ensuring enhanced performance in both camera and object motion control.
5 Conclusion
@@
Acknowledgements: We express our gratitude to Yaowei Li for his adaptation of our proposed MotionCtrl to AnimateDiff [8].
Supplementary Material
The supplementary materials provide additional results achieved with our proposed MotionCtrl, along with in-depth analyses. For a more visual understanding, we strongly recommend readers visit our project page for the video results. The structure of the supplementary materials is as follows:
  • More results of MotionCtrl based on LVDM [9].(Section 6)
  • Results of MotionCtrl when extended to AnimateDiff [8] framework. (Section 7)
  • Visualized results and analyses of ablation studies. (Section 8)
  • More discussions about previous related works. (Section 9).
  • Details of evaluation datasets. (Section 10)
6 More Results of MotionCtrl Deployed on LVDM [9]
In this section, we present additional results of our proposed MotionCtrl model deployed on LVDM [9], focusing on camera motion control, object motion control, and combined motion control. Notably, all results are obtained using the same trained MotionCtrl model, without the need for extra fine-tuning for different camera poses or trajectories.
Specifically, Figure 5 illustrates the outcomes of camera motion control of MotionCtrl guided by 8 basic camera poses, including pan up, pan down, pan left, pan right, zoom in, zoom out, anticlockwise rotation, and clockwise rotation. These poses are visualized in Figure 15 (a). This demonstrates the capability of our MotionCtrl model to integrate multiple basic camera motion controls in a unified model, contrasting with the AnimateDiff model [8] which requires a distinct LoRA model [11] for each camera motion.
Figure 6 showcases the results of camera motion control using MotionCtrl, which is guided by relatively complex camera poses. These complex camera poses are distinct from basic camera poses, as they include elements of camera turning or self-rotation within the same camera pose sequence. The results demonstrate that, given a sequence of camera poses, our MotionCtrl can generate natural videos. The content of these videos aligns with the text prompts, and the camera motion corresponds to the provided complex camera poses. For additional results, please visit our project page.
Figure 7 presents the results of object motion control using MotionCtrl, guided by specific trajectories. When given the same trajectories and different text prompts, MotionCtrl can generate videos featuring different objects, but with identical object motion, as demonstrated in the first four samples of Figure 7. Furthermore, when provided with multiple trajectories within the same video, MotionCtrl is capable of controlling the motion of several objects simultaneously in the same generated video, as shown in the last samples of Figure 7.
Figure 8 provides the results of combining both the camera motion control and object motion control. With the same trajectory but different camera poses, the horse in the generated videos has a different performance.
Figure 5: The results of our proposed MotionCtrl deployed on LVDM [9], guided by 8 basic camera poses: pan up, pan down, pan left, pan right, zoom in, zoom out, anticlockwise rotation, and clockwise rotation (The visualization of these camera poses can be seen in Figure 15 (a)). It’s important to note that all results are achieved using the same MotionCtrl model, without the need for extra fine-tuning for different camera poses.
Figure 6: The results of our proposed MotionCtrl deployed on LVDM [9], guided by relatively complex camera poses. Unlike basic camera poses, which only involve simple directional movements, these complex camera poses incorporate elements of camera turning or self-rotation within the same camera pose sequence. The camera motion in the generated videos closely follows the guided camera poses, while the generated content aligns with the text prompts. For additional results, please visit our project page.
Figure 7: The result of our proposed MotionCtrl deployed on LVDM [9], guided with trajectories. The green points in the trajectories indicate the starting points. Given the same trajectories, our model can generate different objects in accordance with the text prompts, maintaining the same object motion. When multiple trajectories are present in the same video, our model is capable of simultaneously controlling the motion of different objects within the same generated video.
Figure 8: The result of combining camera motion and object motion control of MotionCtrl deployed on LVDM [9]. With the same trajectory but different camera poses, the horse in the generated videos has a different performance.
7 Results of MotionCtrl Deployed on AimateDiff [8]
We also deploy our MotionCtrl on AnimateDiff [8]. Therefore, we can control the motion of the video generated with our fine-tuned AnimateDiff cooperating with various LoRA [11] models in the committee. The results in Figure 11129, and 10 are generated resutls of our MontionCtrl cooperating with different LoRA models provided by in CIVITAI [1]. They demonstrate that our MotionCtrl also works well in camera motion control and object motion control. Related results are also provided on the project page.
Figure 9: The camera motion control results of MotionCtrl deployed on AnimateDiff [8]. They are guided with relatively complex camera poses.
Figure 10: The object motion control results of MotionCtrl deployed on AnimateDiff [8].
Figure 11: The camera motion control results of MotionCtrl deployed on AnimateDiff [8]. They are guided with 8 basic camera poses.
Figure 12: The camera motion control results of MotionCtrl deployed on AnimateDiff [8]. Our MotionCtrl can not only control the camera motion of the generated videos but also their motion speed.
8 More Results and Analyses of Ablation Studies
The manuscript includes numerous ablation studies and detailed analyses. This section provides qualitative results of the ablation studies concerning the integrated position of the Camera Motion Control Module (CMCM) and the training strategy of the Object Motion Control Module (OMCM).
Specifically, Figure 13 demonstrates the generated results of LVDM[9] and our MotionCtrl deployed on LVDM, implemented by integrating CMCM with different components in LVDM. These components include the time embedding block, the cross-attention or self-attention module in the spatial transformers, and temporal transformers (accepted in our final MotionCtrl). The results indicate that MotionCtrl fails to control the camera motion of the generated video, except when integrating CMCM with the temporal transformers. This is primarily because camera motion mainly leads to global and temporal movement in the video. The camera poses integrated into the temporal transformer can effectively fuse the camera movement information into the T2V generation model and yield camera motion.
Figure 14 illustrates the qualitative outcomes of the ablation study ”Dense Trajectories vs. Sparse Trajectories”. The results reveal that the Object Motion Control Module (OMCM) in MotionCtrl, when trained with dense trajectories, fails to control the object motion in the generated video. This is primarily due to the disparity between the dense trajectories in the training phase and sparse trajectories in the inference phase. Furthermore, the model trained on dense trajectories, followed by a fine-tuning phase on sparse trajectories, demonstrates superior precision in object motion control compared to the model trained exclusively on sparse trajectories. This is largely because sparse trajectories are too scattered to be effectively learned, and a model with dense trajectories can provide a more optimal starting point for the learning of sparse trajectories.
Figure 13: The qualitative results of ablation study regarding the integrated position of the Camera Motion Control Module (CMCM) with LVDM [9]. Integrating CMCM of MotionCtrl with the temporal transformers in LVDM significantly improves camera motion control compared to other setups.
Figure 14: The qualitative results of ablation study ”Dense Trajectories v.s. Sparse Trajectories”. The model trained with dense trajectories fails to control the object motion in the generated video. Conversely, the model trained on dense trajectories, followed by fine-tuning on sparse trajectories, exhibits superior precision in object motion control compared to the model trained solely on sparse trajectories.
9 More Discussions about the Related Works
To further illustrate the advantages of our proposed MotionCtrl, we’ve conducted a comparative analysis with previous related works. The comparisons are detailed in Table 4. Models such as AnimateDiff[8] (refers to the motion control LoRA models provided by AnimateDiff), Tune-a-video[26], LAMP[27], and MotionDirector[30] implement motion control by extracting motion from one or multiple template videos. This approach necessitates the training of distinct models for each template video or template video set. Moreover, the motions these methods learned are solely determined by the template video(s), and they fail to differentiate between camera motion and object motion. Similarly, MotionDirector[30] and VideoComposer[25], despite achieving motion control with a unified model guided by motion vectors and trajectories respectively, do not distinguish between camera motion and object motion. In contrast, our proposed MotionCtrl, utilizing a unified model, can independently and flexibly control a wide range of camera and object motions in the generated videos. This is achieved by guiding the model with camera poses and trajectories respectively, offering a more fine-grained control over the video generation process.
AnimateDiff [8] template videos
Tune-a-video [26] template video
LAMP [27] template videos
MotionDirector [30] template videos
VideoComposer [25] motion vectors
DragNUWA [28] trajectories
MotionCtrl (Ours) camera poses & trajectories
Table 4: Differences between our proposed MotionCtrl and related works. Unlike AnimateDiff [8] (which refers to the motion control LoRA model provided by AnimateDiff), Tune-a-video [26], LAMP [27], and MotionDirector [30] that implement motion control by extracting motion from one or a series of template videos and require different models for different template videos, our proposed MotionCtrl uses a unified model. Besides, the motions learned by these methods are determined by the template video(s) and they do not distinguish between camera motion and object motion. On the other hand, although MotionDirector [30] and VideoComposer [25] achieve motion control with a unified model guided by motion vectors and trajectories, respectively, they also do not distinguish between camera motion and object motion. In contrast, our proposed MotionCtrl, with a unified model, can independently and flexibly control the camera motion and object motion of the generated video, guided by camera poses and trajectories, respectively.
10 Details of Evaluation Datasets
@@
In this paper, we construct two evaluation datasets to independently evaluate the efficacy of our proposed MotionCtrl on camera motion control and object motion control.
Camera motion control evaluation dataset. This dataset contains a total of 280 samples. It consists of 8 basic camera poses, 20 complex camera poses, and 10 text prompts. The basic camera poses include pan left, pan right, pan up, pan down, zoom in, zoom out, anticlockwise rotation, and clockwise rotation. The complex camera poses are derived from the testset of Realestate10K [33]. Visualizations of these poses are provided in Figure 15, demonstrating the dataset’s diverse range of camera poses. Besides, The text prompts are listed below:
  • A train travels along a railway through a valley.
  • Rocky coastline with crashing waves.
  • Effiel Tower in Paris.
  • City of Venice, with buildings, river, and boats.
  • A pyramid in the desert.
  • A castle in the forest.
  • A villa in a garden.
  • A fish is swimming in the aquarium tank.
  • A zebra next to a river.
  • A massive, multi-tiered elven palace adorned with flowing waterfalls, its cascades forming staircases between ethereal realms.
Object Motion Control Evaluation Dataset. This evaluation dataset consists of 19 diverse trajectories and 12 text prompts. The trajectories, which include straight lines, curves, shaking lines, and combinations thereof, are illustrated in Figure 16. The text prompts are provided below:
  • A man skateboarding
  • A man surfing.
  • A girl skiing.
  • The rose swaying in the wind.
  • The sunflower swaying in the wind.
  • Two zebras walking on the grass.
  • Two horses walking on the grass.
  • A train passing by the valley.
  • A car running on Mars.
  • A horse running on Mars.
  • A cute dog sitting on the green grass.
  • A feather floating in the air.
@@
(b) 20 Complex Camera Poses from Realestate10K Testset
@@
Figure 15: The Camera Motion Control Evaluation Dataset consists of 8 basic camera poses and 20 complex camera poses, with the complex poses being derived from the test set of Realestate10K. This dataset is utilized to quantitatively assess the effectiveness of our proposed MotionCtrl in controlling a wide range of diverse camera motions in videos generated.
@@
Figure 16: The Object Motion Control Evaluation Dataset encompasses 19 trajectories, where the green and blue points respectively represent the starting and ending points of each trajectory. This dataset is used to quantitatively evaluate the effectiveness of the proposed MotionCtrl in controlling object movements in videos generated.

### v2

@@ 4.3. Ablation Studies
4.3.3. Training Strategy.
Given the limitations of the available training dataset, we propose a multi-step training strategy for MotionCtrl, starting with the CMCM using Realestate10K (Zhou et al., 2018), followed by the OMCM with synthesized object movement trajectories. To thoroughly assess our approach, we experiment with reversing the order and training OMCM before CMCM. This sequence does not impact camera motion control, as OMCM components do not participate in CMCM training. However, it leads to a decrease in object motion control performance since the subsequent training of CMCM adjusts parts of LVDM’s temporal transformers, disrupting the object motion control adaptation achieved during OMCM’s initial training. Thus, our multi-step strategy, though a compromise due to dataset constraints, is deliberately structured to train CMCM before OMCM, ensuring enhanced performance in both camera and object motion control.
4.4. Deploy MotionCtrl on AnimateDiff
We also deploy our MotionCtrl on AnimateDiff (Guo et al., 2023). Therefore, we can control the motion of the video generated with our adjusted AnimateDiff cooperating with various LoRA (Hu et al., 2021) models in the committee. Visualized results of complex camera motion control and object motion control are in Fig. 9 and Fig. 10. More results are in the supplementary materials.
5. Limitations
As an initial exploration into controlling camera and object motion within a unified video generation model, MotionCtrl has demonstrated promising and insightful results. However, controlling the camera and object motion in the same video with both complex camera and complex object trajectories requires a careful design of these trajectories to achieve a natural and harmonious outcome, and the success rate is relatively low. Further research is needed to enhance the accuracy of simultaneously controlling camera and object motion in generated videos.
6. Conclusion
@@
Appendix A Details of Training Data Construction
Augmented-RealEstate10K. The camera motion control module (CMCM) in MotionCtrl is trained with data augmented from RealEstate10K (Zhou et al., 2018). RealEstate10K originally contains videos with annotations of camera poses. To adapt it to our MotionCtrl, we further synthesize captions for each video with Blip2 (Li et al., 2023), an image captioning algorithm. Specifically, we extract frames at specific intervals—the first, quarter, half, three-quarters, and final frames of a video. We then use Blip2 to predict their captions. These captions are concatenated to form a comprehensive description for each video clip. With these captions in place, we train the CMCM on RealEstate10K, enabling effective camera motion control in video generation models such as LVDM (He et al., 2022).
Augmented-WebVid. The object motion control module (OMCM) in MotionCtrl is trained with data augmented from WebVid (Bain et al., 2021). WebVid is a large-scale video dataset equipped with captions and commonly used in the T2V generation task. To adapt it to our MotionCtrl, we further synthesize the object movement trajectories for the videos in WebVid with ParticleSfM (Zhao et al., 2022). Although ParticleSfM is a structure-from-motion system primarily, it incorporates a trajectory-based motion segmentation module utilized for filtering out dynamic trajectories that affect the production of camera trajectories in a dynamic scene. The dynamic trajectories attained by the motion segmentation module exactly fulfill the requirements of our MotionCtrl and we employ this module to synthesize moving object trajectories required by our MotionCtrl. However, despite its effectiveness, ParticleSfM is not time-efficient, requiring approximately 2 minutes to process a 32-frame video. To mitigate the issue of time efficiency, we randomly select 32 frames from each WebVid video, with a frame skip interval $s\in[1,16]$, to synthesize the object movement trajectories. This approach yields a total of 243,000 video clips that fulfill the training requirements for the OMCM.
Appendix B Details of Evaluation Datasets
@@
In this paper, we construct two evaluation datasets to independently evaluate the efficacy of our proposed MotionCtrl on camera and object motion control, respectively.
Camera Motion Control Evaluation Dataset. This dataset contains a total of 407 samples covering two types of camera poses:
  1. (1)
    80 ($8\times 10$) samples constructed with 8 basic camera pose sequences (pan left, pan right, pan up, pan down, zoom in, zoom out, anticlockwise rotation, and clockwise rotation) and 10 prompts.
  2. (2)
    200 ($20\times 10$) samples constructed with 20 relatively complex camera pose sequences randomly selected from the test set of RealEstate10K (Zhou et al., 2018) and 10 prompts.
  3. (3)
    100 samples constructed with 100 relatively complex camera poses of WebVid (Bain et al., 2021) synthesized with ParticleSfM (Zhao et al., 2022) and 100 prompts from VBench (Huang et al., 2024).
  4. (4)
    27 samples constructed with 27 relatively complex camera poses of HD-VILA (Xue et al., 2022) synthesized with ParticleSfM and 27 prompts from VBench (Huang et al., 2024).
To provide an intuitive perception of the camera movement, we visualized the 8 basic camera poses and 20 relatively complex camera poses from RealEstate10K (Zhou et al., 2018) in Fig. 11. As described in the manuscript, the term ”complex camera poses” as used in this work denotes camera movements beyond the basic camera poses list, encompassing camera turning and self-rotation within the same camera pose.
Object Motion Control Evaluation Dataset. This evaluation dataset contains a total of 283 samples constructed with 74 diverse trajectories and 77 prompts. It should be noted that to verify the effectiveness of MotionCtrl in object motion control, our evaluation dataset pairs one trajectory with several different prompts or one prompt with several different trajectories. To provide an intuitive perception of the handcrafted trajectories, 19 trajectories adopted in the evaluation dataset are depicted in Fig. 12.
These evaluation datasets will be released.
@@
(b) 20 Relatively Complex Camera Poses from RealEstate10K Testset
@@
Figure 11. The Camera Motion Control Evaluation Dataset consists of 8 basic camera poses and 20 relatively complex camera poses, with the relatively complex poses being derived from the test set of RealEstate10K. This dataset is utilized to quantitatively assess the effectiveness of our proposed MotionCtrl in controlling a wide range of diverse camera motions in videos generated.
@@
Figure 12. The Object Motion Control Evaluation Dataset encompasses 19 trajectories, where the green and blue points respectively represent the starting and ending points of each trajectory. This dataset is used to quantitatively evaluate the effectiveness of the proposed MotionCtrl in controlling object movements in videos generated.
Appendix C More Quantitative and Qualitative Results
C.1. More Quantitative Results
More Quantitative Comparisons on Relatively Complex Camera Motion Control. In the manuscript, the quantitative results of relatively complex camera poses are statistics from all the complex camera poses sourced from RealEstate10K (Zhou et al., 2018), WebVid (Bain et al., 2021), and HD-VILA (Xue et al., 2022). The statistical results for each dataset are presented in Table 4, demonstrating that our MotionCtrl outperforms VideoComposer (Wang et al., 2023) in both the camera poses extracted from RealEstate10K and those synthesized with ParticleSfM (Zhao et al., 2022) (camera poses of WebVid (Bain et al., 2021) and HD-VILA (Xue et al., 2022)) in terms of camera motion control, text similarity, and generated quality.
Table 4. Quantitative Comparisons with VideoComposer (Wang et al., 2023). Our MotionCtrl performs better in all three sets of relatively complex camera poses from RealEstate10K (Zhou et al., 2018), WebVid (Bain et al., 2021), and HD-VILA (Xue et al., 2022).
RealEstate10K WebVid HD-VILA
Method VideoComposer MotionCtrl VideoComposer MotionCtrl VideoComposer MotionCtrl
CamMC $\downarrow$ 0.1073 0.0840 0.0702 0.0589 0.0953 0.0499
CLIPSIM $\uparrow$ 0.2219 0.2324 0.2147 0.2268 0.2429 0.2473
FID $\downarrow$ 134.97 130.29 106.89 102.13 190.54 159.52
FVD $\downarrow$ 1045.82 934.37 733.09 612.84 1709.59 1129.40
User Study. For a more comprehensive evaluation, we conduct a user study involving 34 participants to assess the results of VideoComposer (Wang et al., 2023) and MotionCtrl. The results were generated using object trajectories and relatively complex camera poses covering datasets from RealEstate10K (Zhou et al., 2018), WebVid (Bain et al., 2021), and HD-VILA (Xue et al., 2022). The assessment included criteria such as Video Quality, Text Similarity, and Motion Similarity. Participants are also asked to express their overall preference for each compared pair. The statistical results in Table 5 demonstrate that over 90 percent of participants preferred our results in all assessment aspects. Although VideoComposer exhibited good performance in motion control conditioned on motion vectors, its generated videos often appeared unnatural and strange due to the object shapes captured by the motion vectors from the reference video. Consequently, users showed a stronger preference for our relatively natural results.
Table 5. User Study. Compared to the results generated with VideoComposer (Wang et al., 2023), our MotionCtrl achieved more preference in all assessment aspect.
Method VideoComposer MotionCtrl
Quality $\uparrow$ 0.0628 0.9372
TextSimilarity $\uparrow$ 0.0772 0.9228
MotionSimilarity $\uparrow$ 0.086 0.9140
OverallPreference $\uparrow$ 0.0739 0.9261
C.2. More Qualitative Results
More Qualitative Comparisons with VideoComposer. We present additional qualitative results comparing VideoComposer (Wang et al., 2023) and our proposed MotionCtrl on relatively complex camera and object trajectories in Fig. 13 and Fig. 14, respectively. These results suggest that MotionCtrl outperforms VideoComposer in both camera and object motion control in generated videos. Furthermore, MotionCtrl’s generated videos exhibit higher quality and its generated content is better aligned with the prompts.
Figure 13. More Qualitative Comparisons with VideoComposer (Wang et al., 2023) on Camera Motion Control. The generated videos of MotionCtrl can better follow the camera poses, whether from RealEstate10K (Zhou et al., 2018) or those synthesized with ParticleSfM (Zhao et al., 2022) on videos of WebVid (Bain et al., 2021) and HD-VILA (Xue et al., 2022). Moreover, the results achieved with MotionCtrl exhibit higher quality.
Figure 14. More Qualitative Comparisons with VideoComposer (Wang et al., 2023) on Object Motion Control. The generated videos of MotionCtrl exhibit a superior ability to follow the trajectories in each frame and achieve higher overall quality.
More of MotionCtrl. In this section, we present additional results of MotionCtrl, focusing on camera motion control, object motion control, and combined motion control. Notably, all results are obtained using the same trained MotionCtrl model, without the need for extra fine-tuning for different camera poses or trajectories.
Specifically, Fig. 17 illustrates the outcomes of camera motion control of MotionCtrl guided by 8 basic camera poses, including pan up, pan down, pan left, pan right, zoom in, zoom out, anticlockwise rotation, and clockwise rotation. These poses are visualized in Fig. 11 (a). This demonstrates the capability of our MotionCtrl model to integrate multiple basic camera motion controls in a unified model, contrasting with the AnimateDiff model (Guo et al., 2023) which requires a distinct LoRA model (Hu et al., 2021) for each camera motion.
Fig. 15 showcases the results of camera motion control using MotionCtrl, which is guided by relatively complex camera poses. These complex camera poses are distinct from basic camera poses, as they include elements of camera turning or self-rotation within the same camera pose sequence. The results demonstrate that, given a sequence of camera poses, our MotionCtrl can generate natural videos. The content of these videos aligns with the text prompts, and the camera motion corresponds to the provided complex camera poses.
Fig. 18 presents the results of object motion control using MotionCtrl, guided by specific trajectories. When given the same trajectories and different text prompts, MotionCtrl can generate videos featuring different objects, but with identical object motion.
Fig. 16 provides the results of combining both the camera motion control and object motion control. With the same trajectory but different camera poses, the horse in the generated videos has a different performance.
Figure 15. The results of our proposed MotionCtrl deployed on LVDM (He et al., 2022), guided by relatively complex camera poses. Unlike basic camera poses, which only involve simple directional movements, these complex camera poses incorporate elements of camera turning or self-rotation within the same camera pose sequence. The camera motion in the generated videos closely follows the guided camera poses, while the generated content aligns with the text prompts.
Figure 16. The result of combining camera motion and object motion control of MotionCtrl deployed on LVDM (He et al., 2022). With the same trajectory but different camera poses, the horse in the generated videos has a different performance.
Table 6. Differences between our proposed MotionCtrl and related works. Unlike AnimateDiff (Guo et al., 2023) (which refers to the motion control LoRA model provided by AnimateDiff), Tune-a-video (Wu et al., 2023b), LAMP (Wu et al., 2023a), and MotionDirector (Zhao et al., 2023) that implement motion control by extracting motion from one or a series of template videos and require different models for different template videos, our proposed MotionCtrl uses a unified model. Besides, the motions learned by these methods are determined by the template video(s) and they do not distinguish between camera motion and object motion. On the other hand, although MotionDirector (Zhao et al., 2023) and VideoComposer (Wang et al., 2023) achieve motion control with a unified model guided by motion vectors and trajectories, respectively, they also do not distinguish between camera motion and object motion. In contrast, our proposed MotionCtrl, with a unified model, can independently and flexibly control the camera motion and object motion of the generated video, guided by camera poses and trajectories, respectively.
AnimateDiff (Guo et al., 2023) template videos
Tune-a-video (Wu et al., 2023b) template video
LAMP (Wu et al., 2023a) template videos
MotionDirector (Zhao et al., 2023) template videos
VideoComposer (Wang et al., 2023) motion vectors
DragNUWA (Yin et al., 2023a) trajectories
MotionCtrl (Ours) camera poses & trajectories
Figure 17. The results of our proposed MotionCtrl deployed on LVDM (He et al., 2022), guided by 8 basic camera poses: pan up, pan down, pan left, pan right, zoom in, zoom out, anticlockwise rotation, and clockwise rotation (The visualization of these camera poses can be seen in Fig. 11 (a)). It’s important to note that all results are achieved using the same MotionCtrl model, without the need for extra fine-tuning for different camera poses.
Figure 18. The result of our proposed MotionCtrl deployed on LVDM (He et al., 2022), guided with trajectories. The green points in the trajectories indicate the starting points. Given the same trajectories, our model can generate different objects in accordance with the text prompts, maintaining the same object motion. When multiple trajectories are present in the same video, our model is capable of simultaneously controlling the motion of different objects within the same generated video.
Appendix D More Results of MotionCtrl Deployed on AimateDiff (Guo et al., 2023)
We also deploy our MotionCtrl on AnimateDiff (Guo et al., 2023). Therefore, we can control the motion of the video generated with our fine-tuned AnimateDiff cooperating with various LoRA (Hu et al., 2021) models in the committee. Results of relatively complex camera motion control and object motion control are in the manuscripts and we provide the results of basic camera motion control here: Fig. 19 and Fig. 20. These results are generated with our MontionCtrl cooperating with different LoRA models provided by in CIVITAI (civitai). They demonstrate that our the generalization of MotionCtrl that can be adapted to different video generation models.
Figure 19. The camera motion control results of MotionCtrl deployed on AnimateDiff (Guo et al., 2023). They are guided with 8 basic camera poses.
Figure 20. The camera motion control results of MotionCtrl deployed on AnimateDiff (Guo et al., 2023). Our MotionCtrl can not only control the camera motion of the generated videos but also their motion speed.
Appendix E More Discussions about the Related Works
To further illustrate the advantages of our proposed MotionCtrl, we’ve conducted a comparative analysis with previous related works. The comparisons are detailed in Table 6. Models such as AnimateDiff(Guo et al., 2023) (refers to the motion control LoRA models provided by AnimateDiff), Tune-a-video(Wu et al., 2023b), LAMP(Wu et al., 2023a), and MotionDirector(Zhao et al., 2023) implement motion control by extracting motion from one or multiple template videos. This approach necessitates the training of distinct models for each template video or template video set. Moreover, the motions these methods learned are solely determined by the template video(s), and they fail to differentiate between camera motion and object motion. Similarly, MotionDirector(Zhao et al., 2023) and VideoComposer(Wang et al., 2023), despite achieving motion control with a unified model guided by motion vectors and trajectories respectively, do not distinguish between camera motion and object motion. In contrast, our proposed MotionCtrl, utilizing a unified model, can independently and flexibly control a wide range of camera and object motions in the generated videos. This is achieved by guiding the model with camera poses and trajectories respectively, offering a more fine-grained control over the video generation process.